Two cars start from points 420 miles apart and travel toward each other. They meet after 3.5 hours. Find the average speed of each car if one travels 30 mph slower than the other.
Answer by richwmiller(17219)

r*t=d
3.5r+3.5*(r+30)=420
r=45
r+30=75